### v1

@@ 6 Conclusion
In this paper, we introduce a new video dataset, named Sekai, for video generation-based world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos from 101 countries and over 750 cities. We develop an efficient and effective toolchain to pre-process and annotate videos. For each video, we annotate location, scene type, weather, crowd density, captions, and camera trajectories. Besides, we introduce a video sampling module to sample the best of the best videos according to the model training budget. Experiments demonstrate the effectiveness and the quality of the dataset. We believe Sekai will benefit the area of video world generation and raise valuable applications.
7 Limitation
Insufficient training. Due to the limited computational resources, we use only a small proportion of Sekai-Real-HQ in model training.
Insufficient camera trajectory annotation. Due to our limited computational resources, for Sekai-Real, we annotate camera trajectories only for partial data.

### v3

@@ 6 Conclusion
In this paper, we have introduced a new video dataset, Sekai, for video generation–based world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos collected from 101 countries and more than 750 cities. We have developed an efficient and effective pipeline to process, filter and annotate the videos. For each video, we annotate location, scene type, weather, crowd density, captions, and camera trajectories. In addition, we present a video sampling module that selects top-tier videos according to the model training budget. Our comprehensive analyses and experiments validate the dataset’s scale, diversity, annotation quality, and effectiveness in supporting world exploration video generation model training. We believe that Sekai will benefit the field of video world generation and inspire valuable future applications.
Acknowledgments This work was supported by the ational Key R&D Program of China No.2022ZD0160101, Natural Science Foundation of China (NSFC) under No. 62172041 and No. 62176021, Shenzhen Science and Technology Program under Grant No. JCYJ20241202130548062, and Natural Science Foundation of Shenzhen under Grant No. JCYJ20230807142703006.